George A. Madrid 31 October 1995

Critique of “Death Watch”

As a big fan of Edgar Allen Poe, elements of this story reminded me of “The Tell-tale Heart.” Perhaps that has affected my expectations, but I thought that the death watch appeared too fast and without strong support from the surrounding narrative. He did try. “...the rain growing stronger, pounding out a staccato rhythm on the glass.”(1.8) and “...the repetitive clicks of her roommate’s iron lung...”(1.5) Unfortunately, it doesn’t hold together, and when the clock does appear it somehow feels wrong.

Also, the descriptions of the clock seem confused. Initially, the clock is described as ticking, so I am very surprised when it is a digital clock. Especially when it only advances every minute. The clocks following night around are described with spinning hands, too.

The style is poetic, and at times, too much so, getting in the way of comprehension. “...losing fingernails on the steep slope of the hole that Cancer dug.”(1.3) The metaphoric descriptions are often mixed. “To Jason, those days were like a cruise...”(1.4) So were those days a cruise or a hole?

David seems to be trying for a dark mood. He needs to be careful not to lighten it with attempts to be cute. The “novice acupuncturist”(2.2) comment is a good example of this problem. Calling the fat man “Pimple”(4.4) also makes me giggle a little too much. When the story is this short, every word must contribute to the mood. The eight minutes before Pimple dies pass in only two lines. More could have been done here to accentuate the mood.

The last paragraph (or two) annoy me. The repetition of the words “That night” no less than six times, probably trying to suggest the ticking of the death watch, only served to distract me and make me look for the real content at the end.

I think more characterization is needed to fill in the story and make me interested. Suicide stories are often difficult since there is such a strong need to make the reader care before the final act. This one is especially challenging since his suicide is self-fulfilling. Perhaps further development between Jason and his wife would help.